# Design Guidance Towards Addressing Over-Reliance on AI in Sensemaking

### 09 Mar 2026 v1

@@ 1. Extended Abstract
The third consideration concerns interaction techniques that enable groups to investigate GenAI-generated awareness information rather than passively receive it. Even when GenAI assessments are presented as secondary encodings, groups need affordances to examine the evidence underlying them in order to engage in autonomous elaboration and discussion rather than simply accepting the differences revealed. Figure 2 shows a hover-for-details interaction: hovering over a lightly shaded axis reveals GenAI’s assessed understanding level, a confidence score, and example discussion quotes that served as evidence (Dülger et al., 2025; Hu et al., 2025; Chen et al., 2021; Farrokhnia et al., 2025). A group observing a discrepancy can read the excerpts, judge whether they agree with the interpretation, and decide whether it warrants further discussion. This positions GenAI outputs as starting points for autonomous elaboration and discussion, supporting the implicit guidance mechanism that GATs are designed to trigger.
@@
These three considerations suggest that integrating GenAI into GATs in a way that provides implicit guidance requires attending not only to what awareness information GenAI generates, but to how it is presented and how groups are supported in engaging with it. Together, they point toward a design space in which GenAI enriches the differences revealed to groups rather than simplifying them into conclusions, preserving and potentially deepening the cognitive conflict that triggers autonomous sensemaking. We bring these considerations to the workshop for discussion: what other design strategies might preserve implicit guidance in AI-supported sensemaking, and how transferable is this mechanism to sensemaking contexts beyond collaborative work and learning?

### 11 Mar 2026 v2

@@ 1. Extended Abstract
The third consideration concerns interaction techniques that enable groups to investigate GenAI-generated awareness information rather than passively receive it. Even when GenAI assessments are presented as secondary encodings, groups need affordances to examine the evidence underlying them in order to engage in autonomous elaboration and discussion rather than simply accepting the differences revealed. Figure 2 shows a hover-for-details interaction: hovering over a lightly shaded axis reveals GenAI’s assessed understanding level, a confidence score, and example discussion quotes that served as evidence (Dülger et al., 2025; Hu et al., 2025a; Chen et al., 2021; Farrokhnia et al., 2025). A group observing a discrepancy can read the excerpts, judge whether they agree with the interpretation, and decide whether it warrants further discussion. This positions GenAI outputs as starting points for autonomous elaboration and discussion, supporting the implicit guidance mechanism that GATs are designed to trigger.
@@
These three considerations suggest that integrating GenAI into GATs in a way that provides implicit guidance requires attending not only to what awareness information GenAI generates, but to how it is presented and how groups are supported in engaging with it. Together, they point toward a design space in which GenAI enriches the differences revealed to groups rather than simplifying them into conclusions, preserving and potentially deepening the cognitive conflict that triggers autonomous sensemaking. Thus, we hope to discuss in this workshop what other design strategies might preserve implicit guidance in AI-supported sensemaking, and how transferable this mechanism is to sensemaking contexts beyond collaborative work and learning. Our prior research contributes expertise spanning machine learning, AI-based systems supporting individual and collaborative knowledge and creative work, interaction design with AI, and cognitive and metacognitive augmentation (Zhao et al., 2026a; Zhao, 2025; Zhao et al., ; Zhao et al., 2024c, b, 2026b, a; Hu et al., 2025b; Lisena et al., 2026; Simperl et al., 2025a, b; Zhang et al., 2025; Zhao, 2021; Luo et al., 2021).